OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = x-9/x+5 and g(x) = -5x-9/x-1 PLEASE HELP I GIVE MEDALS

OpenStudy (anonymous):

@accessdenied PLEASE HELP?

OpenStudy (accessdenied):

It essentially tells you what to do: take f(x) and put it into g(x) and show that it is equal to x, and the same interchanging f(x) and g(x). Your given functions are not well-written here: I cannot tell if your functions are either \( \displaystyle f(x) = x - \frac{9}{x} + 5, \quad g(x) = -5x - \frac{9}{x} - 1\) or \( \displaystyle f(x) = \frac{x - 9}{x + 5}, \quad g(x) = \frac{-5x - 9}{x - 1} \) In either case, however, the process is the same.